# XUE: The CO2-rich terrestrial planet-forming region of an externally irradiated Herbig disk

### v1

@@ 3.2 Stellar host properties
We derived the stellar parameters of XUE 10 by comparing the observed FORS2 spectrum to a grid of PHOENIX stellar atmosphere models (Husser et al., 2013)222http://phoenix.astro.physik.uni-goettingen.de, as described in Appendix C. We used a minimized chi-square test to fit , log10(g), and [Fe/H], and benchmarked the statistical errors on the fitted quantities with a Markov Chain Monte Carlo (MCMC) simulation (see the corner plot in Figure 12). The derived stellar parameters are summarized in Table 1. In particular, the [Fe/H] solutions could be degenerate, having the highest peak centered at 0, and another peak at [Fe/H] 0.5 (see top panel in Figure 12). The near-solar metallicity value of XUE 10 is in line with near-infrared observations of protostars in other star-forming regions, e.g., Orion and Lupus (Biazzo et al., 2011, 2017). Given the derived effective temperature, XUE 10 corresponds to a PMS with spectral type F0–F1 V (Gray & Corbally, 2009). We constrained the visual extinction to an V 4.8 using the Virtual Observatory SED Analyzer (VOSA) tool (Bayo et al., 2008)333http://svo2.cab.inta-csic.es/theory/vosa/ and the archival optical and infrared photometric data on the source (see also Appendix C). From the measurement of the VVV (Variables in the Vía Láctea) VISTA (Visible and Infrared Survey Telescope for Astronomy) apparent J-band magnitude J = 12.004 0.001, we derived a bolometric luminosity of bol = 69 5 ⊙. For this purpose, we adopted the bolometric correction in J-band for a F1 PMS from Pecaut & Mamajek (2013), the constrained V (with an uncertainty of 0.2 from Gaia), and a total-to-selective extinction factor V = 3.3, as measured in NGC 6357 (Massi et al., 2015; Russeil et al., 2017; Fouesneau et al., 2022). The derived luminosity places XUE 10 above the main sequence, consistent with a PMS nature, as expected for the young age of NGC 6357. The measured Gaia DR3 parallax also places XUE 10 in the cluster (Kuhn et al., 2019), suggesting that it has the same age as the other members. Assuming a PARSEC (Bressan et al., 2012) isochrone of 1 Myr, which takes into account extinction, XUE 10 is consistent with a PMS of stellar mass ⋆ = 2.5–3.0 ☉.
@@
In this section we characterize the chemical composition and the main molecular physical parameters, i.e., total column density (tot), gas temperature (gas), and equivalent emitting radius (em) of the detected molecules.
Prior to modeling molecular features in the MIRI spectrum, we subtracted the local dust continuum as described in Appendix A. In the continuum-subtracted spectrum, we employed 0D gas slab radiative transfer models computed assuming LTE gas (Tabone et al., 2023).
@@ 4.1 CO2
We computed the slab models for 12CO2, 13CO2, 16O12C18O, and 16O12C17O using all identified emission lines (blended) and bands per species listed in Table 4. We built the model grid varying gas from 100 to 1200 K, log10(tot) from 15 to 22, and em from 0.01 au to 10 au (in log10-space). We first fitted the 12CO2 species, selecting the wavelength ranges where only 12CO2 is predominant, and derived the best fitting values for tot, gas and em. Then we subtracted its best fitting model, and proceeded with the fit of 13CO2, then 16O12C18O, and lastly 16O12C17O in the same fashion. We assumed that they all emit from the same equivalent emitting area, by fixing the emitting radius to that derived from the 12CO2 fit (Rem = 1.15 au; ?Fixed emitting radius? in Table 2). The isotopologues of 12CO2 are expected to be less abundant and more optically thin than the main species, as evident for CO (Dartois et al., 2003; van Zadelhoff et al., 2001).
@@
4.2 Anomalous C- and O- isotopic ratios in CO2?
Looking at the column density ratios between the different detected CO2 isotopologues, we can see that they are not consistent with the 12C/13C, 16O/18O, and 16O/17O isotopic ratios observed in the interstellar medium (ISM), if we interpret them as abundance ratios in the disk (see Table 3). In particular, the column densities of 16O12C18O and 16O12C17O seem to be enhanced by a factor 4–6, and of 13CO2 by a factor 10 than ordinary.
@@ 4.2 Anomalous C- and O- isotopic ratios in CO2?
The observed spectrum of each isotopologue is actually a mix of optically thin and optically thick emission lines. In fact, the measured optical depth at the line center (hereafter line opacity) in the slab models peaks in an optically thick regime ( 1) around the fundamental -branch of each isotopologue, with maximum averaged line optical depth of 90 (12CO2), 55 (13CO2), 4 (16O12C18O), and 3 (16O12C17O), corresponding to the strongest -branch transitions. The line opacity then gradually decreases into an optically thin regime ( 1), corresponding to the - and -branch lines and hot -branch bandheads at wavelengths shorter than 14 m and longer than 16 m. We can therefore hypothesize that the bulk of each isotopologue emission is primarily governed either by the characteristics of the optically thick -branches, or by those of the optically thinner features. If the optically thick emission dominates, the CO2 isotopologues must trace roughly the same gas reservoir, and the difference in -branch line optical depth can explain the enhanced column densities reported in the top part of Table 2. To test this hypothesis, we simultaneously fitted 12CO2, 13CO2, 16O12C18O, and 16O12C17O at fixed 12CO2 temperature (365 K) and emitting radius (1.15 au). We varied the column density of 12CO2 in log10–space between 17 and 22, scaling the column density of the isotopes by their ISM ratios (see Table 3). In this case, a single slab model takes into account the contribution of all four species. The resulting fit is shown in Figure 4. The best-fit ( = 2.2) column densities are 2.2 1018 cm-2 (13CO2), 5.3 1017 cm-2 (16O12C18O) and 1.5 1017 cm-2 (16O12C17O). We note that these values are within the derived 1 error bars in Table 2. We can see in Figure 4, however, that while the spectrum of 13CO2 is well reproduced in this case, the observed features of the other species are underestimated, including the optically thinner hot bands. The shape of their -branches is also not well reproduced. This finding suggests that 13CO2 emission is well characterized by its optically thick features, and that it comes from the same gas reservoir as 12CO2. This way, a column density of 1019 cm-2 can be explained by the smaller line optical depth of its -branch in comparison to 12CO2. Instead, the emission of 16O12C18O and 16O12C17O must be dominated by their optically thin features, therefore tracing disk layers of different gas temperature and/or equivalent radius than 12CO2.
@@
In order to evaluate the dependence of tot on the equivalent emitting radius, we independently modeled 13CO2, 16O12C18O, and 16O12C17O leaving em as a free parameter (?Free emitting radius? in Table 2 and Figure 13). As expected from the intrinsic parameter degeneracies of the slab modeling approach, in this case a larger number of models of different parameter combinations can reproduce equally well the spectrum. 13CO2 is consistent with tracing a closer (0.49 au) and hotter (609 K) gas reservoir than 12CO2, but at ISM-like column density (5 1018 cm-2). Therefore, it is not possible to determine whether 12CO2 and 13CO2 are co-spatial, which hinders our ability to draw strong conclusions when comparing the 12C/13C abundance ratio to the ISM value (see Table 3). Emission from 16O12C17O is instead consistent with originating at an equivalent radius of 0.8 au and a gas temperature of 467 K, while 16O12C18O seems to emit from a larger radius of 1.2 au, further out than 12CO2, at a lower temperature of 344 K. These gas temperatures are consistent with those inferred from the slab models with fixed emitting radius, while the column densities remain aligned with non-ISM oxygen isotopic ratios, i.e., 1.6 1018 cm-2 for 16O12C18O, and 0.5 1018 cm-2 for 16O12C17O respectively. Given the overall optically thin emission of these isotopologues, though, the derived column densities (and emitting radii) may represent lower limits. We conclude that 16O12C18O and 16O12C17O trace optically thin layers in the disk not co-spatial with 12CO2, and where the oxygen isotopic composition may be anomalous. However, 0D gas slab models do not allow to discriminate between line optical depth effects and abundance patterns. The most plausible conclusion is that we are affected by line optical depth effects also towards 16O12C18O and 16O12C17O. A 2D disk thermochemical analysis, e.g., using ProDiMo (Kamp et al., 2010; Woitke et al., 2016; Rab et al., 2018; Woitke et al., 2024a) or DALI (e.g., Bruderer et al., 2012; Bruderer, 2013) models may be employed to derive the emitting regions in the disk of the different isotopologues and study their abundance profile in the radial and vertical disk direction. To do so, however, the chemical networks of 16O12C18O and 16O12C17O are needed to be implemented in these frameworks.
@@ 4.3 CO and H2O
We observe faint CO emission at 4.9–5.2 m, with a detection at 2 significance above noise level on the brightest lines (Figure 5). We characterized the emission calculating the slab models varying gas from 100 to 3400 K, and log10(tot) from 14 to 20. The map in Figure 13 points toward a column density 2.5 1015 cm-2 and a gas temperature of about 1850 K within an equivalent emitting radius of 5–10 au (see Table 2). As the CO emission is optically thin, we only constrain a lower limit on the column density and on the emitting radius. The fitted gas temperature is within the range typically observed in T Tauri disks (e.g., Ramírez-Tannus et al., 2023; Schwarz et al., 2024; Temmink et al., 2024b). However, it shall be noted that JWST/MIRI observations only cover the high- transitions of the CO -branch, which tend to trace hotter and less dense gas than the low- transitions (i.e., visible in the near-IR range). As a consequence, fitting the CO emission using only the high- transitions yields a higher excitation temperature than the actual gas kinetic temperature (Temmink et al., 2024b).
@@
From Figure 5 we can notice that the best-fit models predict bright CO lines between 5.05 and 5.15 m, which in the data are undetected or 1 above the noise level. We find no evidence of contamination from stellar spectral features at the wavelengths of the CO lines, with the exception of the transition at 5.1332 m, which may be hampered by the stellar spectrum. Therefore, at the resolution of MIRI MRS we may be observing a blend of CO lines in the = 1–0, = 2–1 and = 3–2 vibrational states, where the 2–1 and 3–2 vibrational transitions are (partially) excited in non-LTE conditions. As a result, our LTE slab model overpredicts the line flux of these transitions, which must not emit at the same temperature of the 1–0 transitions (Ramírez-Tannus et al., 2023; Grant et al., 2023; Temmink et al., 2024b). Follow-up observations at higher sensitivity, including the near-IR coverage of the lower hand of the -branch ladder, coupled with non-LTE modeling are needed to increase the signal-to-noise ratio of the individual CO lines, and deliver unbiased physical parameters of its emission in XUE 10.
@@
Besides the detected CO2 and CO emission, we only tentatively detect H2O. In particular, the brightest water lines expected in the 6–8 m and 18–25 m regions lie at 2 above the noise level (see Figure 6), whereas the 13–18 m wavelength range is dominated by carbon dioxide (see Figure 3). As a result, the brightest water lines near 17 m can only be identified at 1 in the residual spectrum of CO2. We therefore could not perform a reliable modeling of this species in the same fashion as for the others. Instead, we computed fiducial slab models of water (plotted in Figure 3 and Figure 6) with the aim of reproducing the lines tentatively identified in each wavelength range. The fiducial model for the 13–18 m region was calibrated to match the H2O line luminosity measured in the residual spectrum of CO2 between 17.04 and 17.39 m (see inset in Figure 3), assuming the same equivalent emitting radius of CO2 (1.15 au). In this way, we were able to estimate an upper limit on the total column density of water of tot 1018 cm-2.
@@
This model is consistent with the warm (400–600 K) component of water typically detected around the 17 m wavelength region in non-irradiated T Tauri disks (Banzatti et al., 2025). The 6–8 m water lines, however, are better reproduced by a model, at equal temperature (650 K), with lower column density (1017 cm-2) within a smaller radius (0.8 au). At longer wavelengths, between 18–25 m, the water spectrum may be explained by the combination of a cold (350 K) and a warm (650 K) component, consistent with mounting evidence that multiple physical components are required to reproduce the mid-infrared water spectrum (Temmink et al., 2024a). Considering the 2 H2O lines in the 7 m region, the upper limit on the total column density would be even lower (Ntot 1017 cm-2). We note that the quadruplet of water lines expected at 23.80–23.96 m (see bottom panel in Figure 6) does not exhibit the line flux ratio characteristic of a strong cool water component, recently used as a proxy of ongoing ice sublimation near the water snowline (Banzatti et al., 2023; Vlasblom et al., 2025). By comparing the detection of the CO2 -branch at 15 m with the the upper limit on H2O from our fiducial model computed at 17 m, we find a lower limit for the CO2/H2O column density ratio of tot (CO2/H2O) 130. However, the lack of robust water line detections can only be partially explained by overlapping CO2 emission. The dust continuum is not a viable alternative explanation, as optically thick dust would hide both water and CO2. A seemingly low water abundance has previously been reported in other Herbig disks (e.g., Fedele et al., 2011). These observations are in contrast to theoretical expectations (Walsh et al., 2015; Antonellini et al., 2016; Agúndez et al., 2018), which predict a water vapor column density of the order of 1019-20 cm-2 in Herbig disks, comparable to the majority of known T Tauri sources. We will assess the origin of a water-poor infrared spectrum in disks around intermediate-mass stars in the following discussion.

### v2

@@ 3.2 Stellar host properties
We derived the stellar parameters of XUE 10 by comparing the observed FORS2 spectrum to a grid of PHOENIX stellar atmosphere models (Husser et al., 2013),222http://phoenix.astro.physik.uni-goettingen.de as described in Appendix C. We used a minimized chi-square test to fit , log10(g), and [Fe/H], and benchmarked the statistical errors on the fit quantities with a Markov chain Monte Carlo (MCMC) simulation (see the corner plot in Figure 12). The derived stellar parameters are summarized in Table 1. In particular, the [Fe/H] solutions could be degenerate, having the highest peak centered at 0, and another peak at [Fe/H] 0.5 (see top panel in Figure 12). The near-solar metallicity value of XUE 10 is in line with near-infrared observations of protostars in other star-forming regions, e.g., Orion and Lupus (Biazzo et al., 2011, 2017). Given the derived effective temperature, XUE 10 corresponds to a PMS with spectral type F0–F1 V (Gray & Corbally, 2009). We constrained the visual extinction to an V 4.8 using the Virtual Observatory SED Analyzer (VOSA) tool (Bayo et al., 2008)333http://svo2.cab.inta-csic.es/theory/vosa/ and the archival optical and infrared photometric data on the source (see also Appendix C). From the measurement of the VVV (Variables in the Vía Láctea) VISTA (Visible and Infrared Survey Telescope for Astronomy) apparent J-band magnitude J = 12.004 0.001, we derived a bolometric luminosity of bol = 69 5 ⊙. For this purpose, we adopted the bolometric correction in J-band for a F1 PMS from Pecaut & Mamajek (2013), the constrained V (with an uncertainty of 0.2 from Gaia), and a total-to-selective extinction factor V = 3.3, as measured in NGC 6357 (Massi et al., 2015; Russeil et al., 2017; Fouesneau et al., 2022). The derived luminosity places XUE 10 above the main sequence, consistent with a PMS nature, as expected for the young age of NGC 6357. The measured Gaia DR3 parallax also places XUE 10 in the cluster (Kuhn et al., 2019), suggesting that it has the same age as the other members. Assuming a PARSEC (Bressan et al., 2012) isochrone of 1 Myr, which takes into account extinction, XUE 10 is consistent with a PMS of stellar mass ⋆ = 2.5–3.0 ☉.
@@
In this section we characterize the chemical composition and the main molecular physical parameters, i.e., total column density (tot), gas temperature (gas), and equivalent emitting radius (em) of the detected molecules. Prior to modeling molecular features in the MIRI spectrum, we subtracted the local dust continuum as described in Appendix A. In the continuum-subtracted spectrum, we employed 0D gas slab radiative transfer models computed assuming LTE gas (Tabone et al., 2023).
@@ 4.1 CO2
We computed the slab models for 12CO2, 13CO2, 16O12C18O, and 16O12C17O using all identified emission lines (blended) and bands per species listed in Table 4. We built the model grid varying gas from 100 to 1200 K, log10(tot) from 15 to 22, and em from 0.01 au to 10 au (in log10-space). We first fit the 12CO2 species, selecting the wavelength ranges where only 12CO2 is predominant, and derived the best fitting values for tot, gas and em. Then we subtracted its best fitting model, and proceeded with the fit of 13CO2, then 16O12C18O, and lastly 16O12C17O in the same fashion. We assumed that they all emit from the same equivalent emitting area, by fixing the emitting radius to that derived from the 12CO2 fit (Rem = 1.15 au; ?Fixed emitting radius? in Table 2). The isotopologues of 12CO2 are expected to be less abundant and more optically thin than the main species, as evident for CO (Dartois et al., 2003; van Zadelhoff et al., 2001).
@@
4.2 Unusual C- and O- isotopic ratios in CO2
When looking at the column density ratios between the different detected CO2 isotopologues, one can see that they are not consistent with the 12C/13C, 16O/18O, and 16O/17O isotopic ratios observed in the interstellar medium (ISM) if they are interpreted as abundance ratios in the disk (see Table 3). In particular, the column densities of 16O12C18O and 16O12C17O seem to be enhanced by a factor of four to six and that of 13CO2 is enhanced by a factor of more than ten compared to the ordinary values.
@@ 4.2 Unusual C- and O- isotopic ratios in CO2
The observed spectrum of each isotopologue is actually a mix of optically thin and optically thick emission lines. In fact, the measured optical depth at the line center (hereafter line opacity) in the slab models peaks in an optically thick regime ( 1) around the fundamental -branch of each isotopologue, with maximum averaged line optical depth of 90 (12CO2), 55 (13CO2), 4 (16O12C18O), and 3 (16O12C17O), corresponding to the strongest -branch transitions. The line opacity then gradually decreases into an optically thin regime ( 1), corresponding to the - and -branch lines and hot -branch bandheads at wavelengths shorter than 14 m and longer than 16 m. We can therefore hypothesize that the bulk of each isotopologue emission is primarily governed either by the characteristics of the optically thick -branches, or by those of the optically thinner features. If the optically thick emission dominates, the CO2 isotopologues must trace roughly the same gas reservoir, and the difference in -branch line optical depth can explain the enhanced column densities reported in the top part of Table 2. To test this hypothesis, we simultaneously fit 12CO2, 13CO2, 16O12C18O, and 16O12C17O at fixed 12CO2 temperature (365 K) and emitting radius (1.15 au). We varied the column density of 12CO2 in log10–space between 17 and 22, scaling the column density of the isotopes by their ISM ratios (see Table 3). In this case, a single slab model takes into account the contribution of all four species. The resulting fit is shown in Figure 4. The best-fit ( = 2.2) column densities are 2.2 1018 cm-2 (13CO2), 5.3 1017 cm-2 (16O12C18O) and 1.5 1017 cm-2 (16O12C17O). We note that these values are within the derived 1 error bars in Table 2. One can see in Figure 4, however, that while the spectrum of 13CO2 is well reproduced in this case, the observed features of the other species are underestimated, including the optically thinner hot bands. The shape of their -branches is also not well reproduced. This finding suggests that 13CO2 emission is well characterized by its optically thick features, and that it comes from the same gas reservoir as 12CO2. This way, a column density of 1019 cm-2 can be explained by the smaller line optical depth of its -branch in comparison to 12CO2. Instead, the emission of 16O12C18O and 16O12C17O must be dominated by their optically thin features, therefore tracing disk layers of different gas temperature and/or equivalent radius than 12CO2.
@@
In order to evaluate the dependence of tot on the equivalent emitting radius, we independently modeled 13CO2, 16O12C18O, and 16O12C17O leaving em as a free parameter (?Free emitting radius? in Table 2 and Figure 13). As expected from the intrinsic parameter degeneracies of the slab modeling approach, in this case a larger number of models of different parameter combinations can reproduce equally well the spectrum. 13CO2 is consistent with tracing a closer (0.49 au) and hotter (609 K) gas reservoir than 12CO2, but at ISM-like column density (5 1018 cm-2). Therefore, it is not possible to determine whether 12CO2 and 13CO2 are co-spatial, which hinders our ability to draw strong conclusions when comparing the 12C/13C abundance ratio to the ISM value (see Table 3). Emission from 16O12C17O is instead consistent with originating at an equivalent radius of 0.8 au and a gas temperature of 467 K, while 16O12C18O seems to emit from a larger radius of 1.2 au, further out than 12CO2, at a lower temperature of 344 K. These gas temperatures are consistent with those inferred from the slab models with fixed emitting radius, while the column densities remain aligned with non-ISM oxygen isotopic ratios, i.e., 1.6 1018 cm-2 for 16O12C18O, and 0.5 1018 cm-2 for 16O12C17O respectively. Given the overall optically thin emission of these isotopologues, though, the derived column densities (and emitting radii) may represent lower limits. We conclude that 16O12C18O and 16O12C17O trace optically thin layers in the disk not co-spatial with 12CO2, and where the oxygen isotopic composition may be anomalous. However, 0D gas slab models do not allow one to discriminate between line optical depth effects and abundance patterns. The most plausible conclusion is that we are affected by line optical depth effects also toward 16O12C18O and 16O12C17O. A 2D disk thermochemical analysis, e.g., using ProDiMo (Kamp et al., 2010; Woitke et al., 2016; Rab et al., 2018; Woitke et al., 2024a) or DALI (e.g., Bruderer et al., 2012; Bruderer, 2013) models may be employed to derive the emitting regions in the disk of the different isotopologues and study their abundance profile in the radial and vertical disk direction. To do so, however, the chemical networks of 16O12C18O and 16O12C17O are needed to be implemented in these frameworks.
@@ 4.3 CO and H2O
We observe faint CO emission at 4.9–5.2 m, with a detection at 2 significance above noise level on the brightest lines (Figure 5). We characterized the emission calculating the slab models varying gas from 100 to 3400 K, and log10(tot) from 14 to 20. The map in Figure 13 points toward a column density 2.5 1015 cm-2 and a gas temperature of about 1850 K within an equivalent emitting radius of 5–10 au (see Table 2). As the CO emission is optically thin, we only constrain a lower limit on the column density and on the emitting radius. The fit gas temperature is within the range typically observed in T Tauri disks (e.g., Ramírez-Tannus et al., 2023; Schwarz et al., 2024; Temmink et al., 2024b). However, it shall be noted that JWST/MIRI observations only cover the high- transitions of the CO -branch, which tend to trace hotter and less dense gas than the low- transitions (i.e., visible in the near-IR range). As a consequence, fitting the CO emission using only the high- transitions yields a higher excitation temperature than the actual gas kinetic temperature (Temmink et al., 2024b).
@@
From Figure 5 one can notice that the best-fit models predict bright CO lines between 5.05 and 5.15 m, which in the data are undetected or 1 above the noise level. We find no evidence of contamination from stellar spectral features at the wavelengths of the CO lines, with the exception of the transition at 5.1332 m, which may be hampered by the stellar spectrum. Therefore, at the resolution of MIRI MRS we may be observing a blend of CO lines in the = 1–0, = 2–1 and = 3–2 vibrational states, where the 2–1 and 3–2 vibrational transitions are (partially) excited in non-LTE conditions. As a result, our LTE slab model overpredicts the line flux of these transitions, which must not emit at the same temperature of the 1–0 transitions (Ramírez-Tannus et al., 2023; Grant et al., 2023; Temmink et al., 2024b). Follow-up observations at higher sensitivity, including the near-IR coverage of the lower hand of the -branch ladder, coupled with non-LTE modeling are needed to increase the signal-to-noise ratio of the individual CO lines, and deliver unbiased physical parameters of its emission in XUE 10.
@@
Apart from the detected CO2 and CO emission, we only tentatively detected H2O. In particular, the brightest water lines expected in the 6–8 m and 18–25 m regions lie at 2 above the noise level (see Figure 6), whereas the 13–18 m wavelength range is dominated by carbon dioxide (see Figure 3). As a result, the brightest water lines near 17 m can only be identified at 1 in the residual spectrum of CO2. We therefore could not perform a reliable modeling of this species in the same fashion as for the others. Instead, we computed fiducial slab models of water (plotted in Figure 3 and Figure 6) with the aim of reproducing the lines tentatively identified in each wavelength range. The fiducial model for the 13–18 m region was calibrated to match the H2O line luminosity measured in the residual spectrum of CO2 between 17.04 and 17.39 m (see inset in Figure 3), assuming the same equivalent emitting radius of CO2 (1.15 au). In this way, we were able to estimate an upper limit on the total column density of water of tot 1018 cm-2.
@@
This model is consistent with the warm (400–600 K) component of water typically detected around the 17 m wavelength region in non-irradiated T Tauri disks (Banzatti et al., 2025). The 6–8 m water lines, however, are better reproduced by a model, at equal temperature (650 K), with lower column density (1017 cm-2) within a smaller radius (0.8 au). At longer wavelengths, between 18–25 m, the water spectrum may be explained by the combination of a cold (350 K) and a warm (650 K) component, consistent with mounting evidence that multiple physical components are required to reproduce the mid-infrared water spectrum (Temmink et al., 2024a). Considering the 2 H2O lines in the 7 m region, the upper limit on the total column density would be even lower (Ntot 1017 cm-2). We note that the quadruplet of water lines expected at 23.80–23.96 m (see bottom panel in Figure 6) does not exhibit the line flux ratio characteristic of a strong cool water component, recently used as a proxy of ongoing ice sublimation near the water snowline (Banzatti et al., 2023; Vlasblom et al., 2025). By comparing the detection of the CO2 -branch at 15 m with the upper limit on H2O from our fiducial model computed at 17 m, we find a lower limit for the CO2/H2O column density ratio of tot (CO2/H2O) 130. However, the lack of robust water line detections can only be partially explained by overlapping CO2 emission. The dust continuum is not a viable alternative explanation, as optically thick dust would hide both water and CO2. A seemingly low water abundance has previously been reported in other Herbig disks (e.g., Fedele et al., 2011). These observations are in contrast to theoretical expectations (Walsh et al., 2015; Antonellini et al., 2016; Agúndez et al., 2018), which predict a water vapor column density of the order of 1019-20 cm-2 in Herbig disks, comparable to the majority of known T Tauri sources. We assess the origin of a water-poor infrared spectrum in disks around intermediate-mass stars in the following discussion.